Alpha and Beta are divisions within the same company. The managers of both divisions are evaluated based on their own divisioRequired: 1. Refer to case 1 shown above. Alpha Division can avoid $4 per unit in commissions on any sales to Beta Division.3. Refer to case 3 shown above. Assume that Beta Division is now receiving an 3% price discount from the outside supplier. a.
Answer #1

1.a.The Alpha division is operating at full capacity, the lowest acceptable transfer price from the perspective of the Alpha Division = Market Price – Costs Avoided

= $101-$4 = $97

b. highest acceptable transfer price from the perspective of the Beta Division = purchase price being paid to outside supplier

= $94

c.No range of acceptable transfer prices. No, managers will not agree for a transfer

2.a. The Alpha division is operating at full capacity, the lowest acceptable transfer price from the perspective of the Alpha Division = Market Price – Costs Avoided

= $39-$4 = $35          

b. highest acceptable transfer price from the perspective of the Beta Division = purchase price being paid to outside supplier

= $38

c.Range of Acceptable transfer price = $35-$38. Yes, managers will agree for a transfer> yes, there will be disagreement as both the managers will try to maximise their own profits

d.Loss in profits = (38-35)*74000

= $222,000

3.Spare capacity for alpha division

a.the lowest acceptable transfer price from the perspective of the Alpha Division = variable cost per unit

= $38

b. highest acceptable transfer price from the perspective of the Beta Division = purchase price being paid to outside supplier = 63*(1-3%) = $61.11

c.Range = $38-$61.11. Yes, ,managers will agree since benefit for both

d.ROI will increase since the transfer price is greater than variable cost and space capacity exists. Hence, transfer will lead to additional profits.

4.Lowest acceptable transfer price = cost incurred+benefit lost

= 29 + (47-33)/2

= $36 per unit